# Rotationally invariant dynamical lattice regulators for Euclidean quantum field theories

### v1

@@ 2.1 Abstract lattice, fields, and embedding
  • for finite–volume simulations, the discrete torus
@@
In either case we denote by $\hat{\mu}$, $\mu=1,\dots,d$, the unit vectors in the coordinate directions and write $n+\hat{\mu}$ for the nearest neighbour of $n$ in direction $\mu$.
@@
which assigns to each abstract site $n\in\Lambda$ a point $x(n)\in\mathbb{R}^{d}$ in physical Euclidean space. In the torus formulation one instead takes $x(n)\in T^{d}$ and works with the shortest–distance representative of coordinate differences; the local constructions in Sections 2.2–2.6 are unchanged.
@@
In this viewpoint the dynamical–lattice regulator is simply a local lattice field theory with additional scalar degrees of freedom $x_{\mu}(n)$ and a special pattern of couplings. This extended–field formulation will be crucial in Section 4, where reflection positivity is established by letting the Osterwalder–Schrader reflection act on the indices $n$ and treating $x$ on the same footing as the other fields.
@@ 2.2 Local Euclidean geometry and admissible embeddings
We now spell out the local Euclidean geometry induced by an embedding $x:\Lambda\to\mathbb{R}^{d}$ of the abstract lattice and formulate our admissibility conditions. Throughout this subsection we work on $\mathbb{R}^{d}$; the periodic (torus) version will be discussed in Section 2.4.
@@
The components of $g(n)$ are $g_{\mu\nu}=e_{\mu}(n)\cdot e_{\nu}(n)$. Whenever $g(n)$ is positive definite we denote the components of its inverse by $g^{\mu\nu}(n)$.
We restrict attention to embeddings for which each pair $E(n)$ and $E^{\theta}\!(n)$ is uniformly nondegenerate and uniformly shape regular, in a way that is stable under global Euclidean motions. A convenient local formulation is the following.
@@
For $x\in\hat{\mathcal{X}}_{\mathrm{adm}}(a)$ introduce the dimensionless embedding $y(n):=a^{-1}x(n)$. Then in terms of $y$, the admissibility conditions (2.9) become aa–independent bounds, meaning that $\hat{\mathcal{X}}_{\mathrm{adm}}(a)$ is a simple rescaling of $\hat{\mathcal{X}}_{\mathrm{adm}}(1)$.
@@
By scaling, we can assume $a=1$. Fix a site $n$ and consider the set $K$ of all frames $E\in\mathbb{R}^{d\times d}$ satisfying (2.9). By construction, $K$ is a closed and bounded set in the space of invertible matrices, and thus the continuous function $A\mapsto\|A^{-1}\|$ achieves its maximum in $K$. ∎
The content of Lemma 2.3 is that the eigenvalues of the local metric $g(n)$ are uniformly bounded above and below away from 0. In numerical–analysis language this is a standard “shape–regularity” condition: the local affine cell determined by the corner frame $E(n)$ cannot be arbitrarily skinny or flat. In particular, (2.9) excludes local collapses of volume and guarantees that the embedding $x:\Lambda\to\mathbb{R}^{d}$ is everywhere locally an immersion with controlled condition number. Global properties such as non–self–intersection will be enforced by restricting $x$ to a suitable admissible component in Definition 2.5 and, if desired, by adding further constraints as discussed in the next remark.
@@
    Cell shape regularity. If desired, one may impose a standard finite–element shape–regularity assumption on the embedded 22–cells: each quadrilateral image is required to be convex and to have uniformly bounded chunkiness parameter $h_{Q}/\rho_{Q}$ (diameter over inradius), uniformly in the lattice and in the configuration. Equivalently, one may enforce a uniform lower bound on all interior angles together with a uniform bound on edge–length ratios; any of the standard equivalent criteria may be used.
@@
    $|x(n)-x(m)|\geq\delta_{\mathrm{sep}}a>0\qquad\text{for all $n\neq m$}.$ (2.15)
@@
Let $x_{0}(n):=a\,n$ be the regular embedding at scale $a>0$. The principal admissible component $\mathcal{X}_{\mathrm{adm}}(a)$ is the connected component of $\hat{\mathcal{X}}_{\mathrm{adm}}(a)$ containing $x_{0}$, i.e.,
@@
Working in $\mathcal{X}_{\mathrm{adm}}$ ensures that the dynamical mesh remains a controlled distortion of the regular hypercubic grid: shape–regularity is enforced uniformly by Lemma 2.3, and no topological defects (such as changes in the combinatorial structure or global self–intersections) are introduced by the fluctuations of $x$. This sector is the natural setting for both the perturbative Symanzik analysis and the nonperturbative reflection–positivity arguments that follow.
@@ 2.3 Geometry measure and dynamics
where $S_{x}[x]$ is a local, $\mathrm{E}(d)$–invariant action on $\mathcal{X}_{\mathrm{adm}}$ and $Z_{x}$ is the corresponding normalisation factor. The hard admissibility constraint is implemented by restricting the domain of integration to $\mathcal{X}_{\mathrm{adm}}$; we do not insert a separate indicator function into the integrand.
@@
so that $d\mu_{x}$ is just the restriction of the flat product measure $Dx$ to $\mathcal{X}_{\mathrm{adm}}$. The admissibility conditions from Definition 2.1 and Remark 2.4 then appear as hard walls in configuration space: any proposed configuration outside $\mathcal{X}_{\mathrm{adm}}$ is assigned infinite action and zero measure. Even in this minimal setting the geometry does not behave as a collection of non-interacting free variables. Because $\mathcal{X}_{\mathrm{adm}}$ is a nontrivial subset of $(\mathbb{R}^{d})^{\Lambda}$, the restriction $x\in\mathcal{X}_{\mathrm{adm}}$ induces effective interactions between neighbouring sites via purely entropic effects: the volume in configuration space available to a site $x(n)$ depends on the positions of its neighbours through the local admissibility constraints. We return to the implications of this “entropic repulsion” in the discussion of Monte–Carlo dynamics.
In $d=2$ our exploratory simulations suggest that the minimal measure $S_{x}^{\min}\equiv 0$ together with the hard walls can already yield a stable, nondegenerate ensemble. In higher dimensions, especially in $d=4$, this is not guaranteed: entropic effects alone may admit phases with long–range geometric correlations or crumpling/roughening phenomena in which the lattice spacing ceases to define a clean regulator scale. For four–dimensional applications we therefore expect that an explicit stiffness term (for instance (2.19) or variants) will be necessary in practice to remain in a short–range geometry regime (SR), that will be formulated below.
@@
Penalties of the form (2.19) are designed to suppress long–wavelength distortions of the local frame and thus to promote a massive (finite–correlation–length) geometry sector once the global $\mathrm{SE}(d)$ zero modes are fixed/quotiented. In the language of Section 4.3, such choices are intended to place the geometry in a regime compatible with the short–range hypothesis (SR).
@@
Universality and the role of $S_{x}$. From the viewpoint of effective field theory, different choices of the geometry action $S_{x}$ define different microscopic regulators, but as long as $S_{x}$ is local, $\mathrm{E}(d)$–invariant, and compatible with the admissibility conditions, they are expected to lie in the same universality class. One may also view the choice of $S_{x}$ (more precisely, the local weight $e^{-S_{x}[x]}$ together with the admissible–set restriction) as selecting an effective submanifold measure on the space of embedded lattices inside the continuum configuration space: it specifies how the enlarged, finite–dimensional family of fields induced by the embedding $x$ is sampled. This reweighting is not required by an exact local redundancy, but changing it (within the same locality and symmetry class) is expected to affect the continuum limit only through symmetry–allowed local counterterms. Integrating out the geometry field induces local, higher–dimensional operators in the effective action for the gauge and matter fields; these operators are constrained by gauge invariance, Euclidean invariance, and reflection positivity, and are therefore irrelevant in the renormalisation–group sense in the continuum limit. We will see an explicit perturbative example of this mechanism in the Symanzik expansion for a scalar theory in §4.4.
@@
Global zero modes versus local twisting. The embedding variables $x(n)$ enjoy an exact global $\mathrm{SE}(d)$ symmetry, $x(n)\mapsto Rx(n)+b$, acting uniformly on all sites. In infinite volume this produces non–normalisable zero modes (and in finite volume it contributes only an irrelevant overall group–volume factor). In practice one therefore fixes/quotients these global modes by a convenient gauge choice, for example by pinning the centre of mass and removing global rotations (e.g. by fixing the orientation of a chosen local frame). This step does not affect local observables and is entirely analogous to removing rigid–motion zero modes in molecular simulations.
@@
The practically relevant question is not whether the regulator is formally $\mathrm{SE}(d)$–covariant (it is, by construction), but how this covariance is realised in the ensemble in a way compatible with universality. Our Symanzik discussion is conditional on the short–range geometry hypothesis (SR): in lattice units the geometry sector must remain massive so that integrating out $x$ produces only local, irrelevant corrections. Within such an (SR) regime, rigid global motions cannot implement $\mathrm{SE}(d)$ covariance at the level of local probes (they carry no local geometric information), and any attempt to enforce isotropy through genuinely long–range geometric modes would violate (SR) and could in principle change the universality class. Thus, in the intended operating regime, $\mathrm{SE}(d)$ covariance that is felt locally must come from the local embedding fluctuations—twisting, shearing, and local reorientation from cell to cell—which average short–distance observables over many local frames while keeping the underlying hypercubic connectivity fixed.
@@
Monte–Carlo dynamics. In numerical simulations the geometry field $x$ is updated by local Metropolis moves of the form $x(n)\to x(n)+\delta x$ with $\delta x$ drawn from a symmetric proposal distribution, followed by an admissibility check at the sites affected by the move and an acceptance test based on the local change in $S_{x}[x]$ and in the field action. Because $S_{x}$ is local, the acceptance probability can be evaluated from a uniformly bounded neighbourhood of $n$, and the resulting Markov chain is manifestly local on the abstract lattice. A detailed description of the update scheme and its ergodicity properties is deferred to the numerics section §5.
@@
Computational cost. From an algorithmic standpoint, the dynamical geometry sector adds only a mild overhead compared to standard lattice gauge theory. A local proposal $x(n)\mapsto x(n)+\delta x$ affects only the finite set of terms in $S_{x}$ and in $S_{\mathrm{fields}}$ that involve $n$ and its nearest neighbours, so both the admissibility check and the Metropolis acceptance ratio can be evaluated in $O(1)$ work per proposal, with a constant depending only on the dimension and the chosen locality range, not on the lattice volume. A full “sweep” of geometry updates therefore costs $O(|\Lambda|)$, just as in standard LGT. In practice the geometry update can be interleaved with the usual gauge and matter updates (or combined in a hybrid scheme), and the method remains compatible with existing locality-based implementations; the principal new ingredient is the additional bookkeeping of the local geometric data needed to evaluate the affected action terms.
@@ 2.4 Torus version and periodic implementation
For most analytical questions (global $\mathrm{E}(d)$ invariance, reflection positivity, Symanzik expansion) it is convenient to regard the dynamical lattice as embedded into the full Euclidean space $\mathbb{R}^{d}$. In numerical simulations, however, one typically works at finite volume with periodic boundary conditions, i.e. on a dd–dimensional discrete torus.
@@
    $\Lambda:=(\mathbb{Z}/N\mathbb{Z})^{d},$ (2.20)
@@
    $x:\Lambda\to\mathbb{T}^{d}:=(\mathbb{R}/L\mathbb{Z})^{d}$ (2.21)
@@
In both formulations the local admissibility conditions are imposed on the edge vectors $e_{\mu}(n)$, and the principal admissible component $\mathcal{X}_{\mathrm{adm}}$ is understood modulo periodicity. Global non–self–intersection is likewise interpreted in the periodic sense: cells may wrap around the torus, but configurations in which large portions of the mesh fold back and concentrate in a small region are excluded by the same local shape–regularity and volume constraints as in the $\mathbb{R}^{d}$ setting.
The passage from $\mathbb{R}^{d}$ to $\mathbb{T}^{d}$ does not affect any of the local arguments in this paper. In particular, the construction of the geometry measure $d\mu_{x}$ and of the field actions remains unchanged, and the analysis of Euclidean invariance and reflection positivity is carried out on $\mathbb{R}^{d}$ and then transferred to the periodic case in the standard way. For the Monte–Carlo implementation we restrict to a finite periodic lattice and update the geometry field subject to these periodic admissibility constraints; the details of the algorithm are deferred to the numerical section.

### v2

@@ 2.1 Abstract lattice, fields, and embedding
  • for finite volume simulations, the discrete torus
@@
In either case we denote by $\hat{\mu}$, for $\mu=1,\dots,d$, the unit vectors in the coordinate directions and write $n+\hat{\mu}$ for the nearest neighbour of $n$ in direction $\mu$.
@@
which assigns to each abstract site $n\in\Lambda$ a point $x(n)\in\mathbb{R}^{d}$ in physical Euclidean space. In the torus formulation one instead takes $x(n)\in\mathbb{T}^{d}$ and works with the shortest distance representative of coordinate differences; the local constructions in what follows are unchanged.
@@
In this viewpoint DLR is simply a local lattice field theory with additional scalar degrees of freedom $x_{\mu}(n)$ and a special pattern of couplings. This extended-field formulation will be crucial in Section 4, where reflection positivity is established by letting the Osterwalder–Schrader reflection act on the indices $n$ and treating $x$ on the same footing as the other fields.
@@ 2.2 Local Euclidean geometry and admissible embeddings
We spell out the local Euclidean geometry induced by an embedding $x:\Lambda\to\mathbb{R}^{d}$ of the abstract lattice and formulate our admissibility conditions. Throughout this subsection we work on $\mathbb{R}^{d}$; the periodic (torus) version will be discussed in Section 2.4.
@@
In $d=2$, $V(n)$ is the (unsigned) area of the outgoing edge parallelogram at $n$; for a non–affine embedded plaquette it need not coincide with the Euclidean area of the quadrilateral cell. The components of $g(n)$ are $g_{\mu\nu}=e_{\mu}(n)\cdot e_{\nu}(n)$. Whenever $g(n)$ is positive definite we denote the components of its inverse by $g^{\mu\nu}(n)$.
We restrict attention to embeddings for which each pair $E(n)$ and $E^{\theta}\!(n)$ is uniformly nondegenerate and uniformly shape regular, in a way that is stable under global proper Euclidean motions. A convenient local formulation is the following.
@@
For $x\in\hat{\mathcal{X}}_{\mathrm{adm}}(a)$ introduce the dimensionless embedding $y(n)=a^{-1}x(n)$. Then in terms of $y$, the admissibility conditions (2.9) become aa–independent bounds, meaning that $\hat{\mathcal{X}}_{\mathrm{adm}}(a)$ is a simple rescaling of $\hat{\mathcal{X}}_{\mathrm{adm}}(1)$.
@@
Fix $n$. Write the rescaled frames $\widetilde{E}=a^{-1}E(n)$ and $\widetilde{E}^{\theta}=a^{-1}E^{\theta}(n)$. Then the defining bounds for admissibility (edge lengths $\sim a$ and oriented volumes $\sim a^{d}$) become uniform bounds for $\widetilde{E}$ and $\widetilde{E}^{\theta}$ at scale $a=1$. The set of all such $\widetilde{E}$ is closed and bounded in the space of invertible matrices with $\det\widetilde{E}$ bounded away from 0, hence compact. The continuous map $A\mapsto\|A^{-1}\|$ attains a finite maximum on this set, say $\|\widetilde{E}^{-1}\|\leq C$ and $\|(\widetilde{E}^{\theta})^{-1}\|\leq C$ with $C$ depending only on $(C_{\ell},c_{V},d)$. Finally, we note
$E(n)^{-1}=(a\widetilde{E})^{-1}=a^{-1}\widetilde{E}^{-1},\qquad(E^{\theta}(n))^{-1}=a^{-1}(\widetilde{E}^{\theta})^{-1},$ (2.12)
establishing the lemma with $C^{*}=2C$. ∎
The content of Lemma 2.3 is that the eigenvalues of the local metrics are uniformly bounded above and below away from 0. In numerical–analysis language this is a standard “shape–regularity” condition: the local affine cells determined by the corner frames cannot be arbitrarily skinny or flat. In particular, (2.9) excludes local collapses of volume and guarantees that the embedding $x:\Lambda\to\mathbb{R}^{d}$ is everywhere locally an immersion with controlled condition number. Global properties such as non–self–intersection will be enforced by restricting $x$ to a suitable admissible component in Definition 2.5 and, if desired, by adding further constraints as discussed in the next remark.
@@
    All-corner volume bounds. One may strengthen the determinant condition in (2.9) by requiring nondegeneracy of all $2^{d}$ corner frames incident to $n$. Concretely, for each site $n$ and each choice of forward/backward directions $\sigma\in\{\pm 1\}^{d}$, let $E^{\sigma}(n)$ be the $d\times d$ “corner frame” obtained by taking, in each coordinate direction $\mu$, the edge from $n$ to $n+\sigma_{\mu}\hat{\mu}$. Then impose the uniform oriented-volume bound
    $\min_{\sigma\in\{\pm 1\}^{d}}\ \Big(\prod_{\mu=1}^{d}\sigma_{\mu}\Big)\,\det E^{\sigma}(n)\ \geq\ c_{V}a^{d}\qquad\text{for all }n\in\Lambda.$ (2.15)
    This rules out fold-overs at any corner and is convenient when actions involve mixed forward/backward stencils in several directions.
  • (A4)
@@
    Cell shape regularity. If desired, one may impose a standard finite–element shape–regularity assumption on the embedded cells: taking the $2d$ case for example, each quadrilateral image is required to be convex and to have uniformly bounded chunkiness parameter $h_{Q}/\rho_{Q}$ (diameter over inradius), uniformly in the lattice and in the configuration. Equivalently, one may enforce a uniform lower bound on all interior angles together with a uniform bound on edge–length ratios; any of the standard equivalent criteria may be used.
@@
    $|x(n)-x(m)|\geq\delta_{\mathrm{sep}}a\qquad\text{for all $n\neq m$}.$ (2.17)
@@
Let $x_{0}(n)=a\,n$ be the regular embedding at scale $a>0$. The principal admissible component $\mathcal{X}_{\mathrm{adm}}(a)$ is the connected component of $\hat{\mathcal{X}}_{\mathrm{adm}}(a)$ containing $x_{0}$, i.e.,
@@
Working in $\mathcal{X}_{\mathrm{adm}}$ ensures that the dynamical mesh remains a controlled distortion of the regular hypercubic grid: shape regularity is enforced uniformly by Lemma 2.3, and no topological defects (such as changes in the combinatorial structure or global self-intersections) are introduced by the fluctuations of $x$. This sector is the natural setting for both the perturbative Symanzik analysis and the nonperturbative reflection positivity arguments that follow.
@@ 2.3 Geometry measure and dynamics
where $S_{x}[x]$ is local and depends on $x$ only through Euclidean invariants of the local edge data (hence is $\mathrm{SE}(d)$–covariant on $\mathcal{X}_{\mathrm{adm}}$), and $Z_{x}$ is the corresponding normalisation factor. The hard admissibility constraint is implemented by restricting the domain of integration to $\mathcal{X}_{\mathrm{adm}}$; we do not insert a separate indicator function into the integrand.
@@
so that $d\mu_{x}$ is just the restriction of the flat product measure $Dx$ to $\mathcal{X}_{\mathrm{adm}}$. The admissibility conditions from Definition 2.1 and Remark 2.4 then appear as hard walls in configuration space: any proposed configuration outside $\mathcal{X}_{\mathrm{adm}}$ is assigned infinite action and zero measure. Even in this minimal setting the geometry does not behave as a collection of non-interacting free variables. Because $\mathcal{X}_{\mathrm{adm}}$ is a nontrivial subset of $(\mathbb{R}^{d})^{\Lambda}$, the restriction $x\in\mathcal{X}_{\mathrm{adm}}$ induces effective interactions between neighbouring sites via purely entropic effects: the volume in configuration space available to a site $x(n)$ depends on the positions of its neighbours through the local admissibility constraints.
In $d=2$ our exploratory simulations suggest that the minimal measure $S_{x}^{\min}\equiv 0$ together with the hard walls can already yield a stable, nondegenerate ensemble. In higher dimensions, especially in $d=4$, this is not guaranteed: entropic effects alone may admit phases with long–range geometric correlations or crumpling/roughening phenomena in which the lattice spacing ceases to define a clean regulator scale. For four–dimensional applications we therefore expect that an explicit stiffness term (for instance (2.21) or variants) will be necessary in practice to remain in a short-range geometry regime (SR), that will be formulated below.
@@
Penalties of the form (2.21) are designed to suppress long–wavelength distortions of the local frame and thus to promote a massive (finite–correlation–length) geometry sector once the global $\mathrm{SE}(d)$ zero modes are fixed/quotiented. In the language of Section 4.3, such choices are intended to place the geometry in a regime compatible with the short-range geometry hypothesis (SR).
@@
Universality and the role of $S_{x}$. From the viewpoint of effective field theory, different choices of the geometry action $S_{x}$ define different microscopic regulators, but as long as $S_{x}$ is local, $\mathrm{E}(d)$–invariant, and compatible with the admissibility conditions, they are expected to lie in the same universality class. One may also view the choice of $S_{x}$ (more precisely, the local weight $e^{-S_{x}[x]}$ together with the admissible–set restriction) as selecting an effective submanifold measure on the space of embedded lattices inside the continuum configuration space: it specifies how the enlarged, finite-dimensional family of fields induced by the embedding $x$ is sampled. This reweighting is not required by an exact local redundancy, but changing it (within the same locality and symmetry class) is expected to affect the continuum limit only through symmetry–allowed local counterterms. Integrating out the geometry field induces local, higher–dimensional operators in the effective action for the gauge and matter fields; these operators are constrained by gauge invariance, Euclidean invariance, and reflection positivity, and are therefore irrelevant in the renormalisation–group sense in the continuum limit. We will see an explicit perturbative example of this mechanism in the Symanzik expansion for a scalar theory in §4.4.
@@
Global zero modes versus local twisting. The embedding variables $x(n)$ enjoy an exact global $\mathrm{SE}(d)$ symmetry, $x(n)\mapsto Rx(n)+b$, acting uniformly on all sites. In infinite volume this produces non–normalisable zero modes, and in finite volume an irrelevant overall group–volume factor. In practice one therefore fixes or quotients these global modes by a convenient convention, e.g., pinning the centre of mass and fixing the orientation of a chosen local frame. This step does not affect local observables and is entirely analogous to removing rigid–motion zero modes in molecular simulations.
@@
The practically relevant question is not whether the regulator is formally $\mathrm{SE}(d)$–covariant (it is, by construction), but how this covariance is realised in the ensemble in a way compatible with universality. Our Symanzik discussion is conditional on the short–range geometry hypothesis (SR); under this assumption, integrating out $x$ produces only local, irrelevant corrections in the effective action. Within this regime, rigid global motions cannot implement $\mathrm{SE}(d)$ covariance at the level of local probes (they carry no local geometric information), and any attempt to enforce isotropy through genuinely long–range geometric modes would violate (SR) and could in principle change the universality class. Thus, in the intended operating regime, $\mathrm{SE}(d)$ covariance that is felt locally must come from the local embedding fluctuations—twisting, shearing, and local reorientation from cell to cell—which average short distance observables over many local frames while keeping the underlying hypercubic connectivity fixed.
@@
Monte Carlo dynamics. In numerical simulations the geometry field $x$ is updated by local Metropolis moves of the form $x(n)\to x(n)+\delta x$ with $\delta x$ drawn from a symmetric proposal distribution, followed by an admissibility check at the sites affected by the move and an acceptance test based on the local change in $S_{x}[x]$ and in the field action. Because $S_{x}$ is local, the acceptance probability can be evaluated from a uniformly bounded neighbourhood of $n$, and the resulting Markov chain is manifestly local on the abstract lattice. A detailed description of the update scheme and its ergodicity properties is deferred to the numerics section §5.
@@
From an algorithmic standpoint, the dynamical geometry sector adds only a mild overhead compared to standard lattice gauge theory. A local proposal $x(n)\mapsto x(n)+\delta x$ affects only the finite set of terms in $S_{x}$ and in $S_{\mathrm{fields}}$ that involve $n$ and its nearest neighbours, so both the admissibility check and the Metropolis acceptance ratio can be evaluated in $O(1)$ work per proposal, with a constant depending only on the dimension and the chosen locality range, not on the lattice volume. A full “sweep” of geometry updates therefore costs $O(|\Lambda|)$, just as in standard LGT. In practice the geometry update can be interleaved with the usual gauge and matter updates (or combined in a hybrid scheme), and the method remains compatible with existing locality-based implementations; the principal new ingredient is the additional bookkeeping of the local geometric data needed to evaluate the affected action terms.
@@ 2.4 Torus version and periodic implementation
For most analytical questions, it is convenient to regard the dynamical lattice as embedded into the full Euclidean space $\mathbb{R}^{d}$. In numerical simulations, however, one typically works at finite volume with periodic boundary conditions, i.e., on a dd–dimensional discrete torus.
@@
    $\Lambda=(\mathbb{Z}/N\mathbb{Z})^{d},$ (2.22)
@@
    $x:\Lambda\to\mathbb{T}^{d}=(\mathbb{R}/L\mathbb{Z})^{d}$ (2.23)
@@
In both formulations the local admissibility conditions are imposed on the edge vectors $e_{\mu}(n)$, and the principal admissible component $\mathcal{X}_{\mathrm{adm}}$ is understood modulo periodicity. Global non–self–intersection is likewise interpreted in the periodic sense: cells may wrap around the torus, but configurations in which large portions of the mesh fold back and concentrate in a small region are excluded by the same local shape regularity and volume constraints as in the $\mathbb{R}^{d}$ setting.
The passage from $\mathbb{R}^{d}$ to $\mathbb{T}^{d}$ does not affect any of the local arguments in this paper. In particular, the construction of the geometry measure $d\mu_{x}$ and of the field actions remains unchanged, and the analysis of Euclidean invariance and reflection positivity is carried out on $\mathbb{R}^{d}$ and then transferred to the periodic case in the standard way. For the Monte Carlo implementation we restrict to a finite periodic lattice and update the geometry field subject to these periodic admissibility constraints; the details of the algorithm are deferred to the numerical section.